# Shilong Liu 5, Haoyu Dong6, Zihao Lin4, Yixin Wang1

### v1

@@ 4 Experimental Settings
MMedAgent is initialized by LLaVA-Med 60K-IM, instruction-tuned using LoRA Hu et al. (2021) for 15 epochs, and conducted over approximately 72 hours on two 80G NVIDIA A100 GPUs. The rank of LoRA is set to 128, and the training batch size is set to 48. We employ AdamW Loshchilov and Hutter (2019) as the optimizer alongside a cosine learning rate schedule peaking at 2e-4. We generate 48K instruction-tuning data, consisting of 15K augmented VQA instruction following the method from LLaVA-Plus Liu et al. (2023a) derived from 60K inline mentions Li et al. (2023), 10K data points for detection, 3K for RAG, 5K each for segmentation, classification, MRG, and G-Seg. Data sources are shown in Table 1.
@@ 5.1.2 Experimental Results
As illustrated in Table 2, MMedAgent significantly outperforms all other baselines on various tasks. Notably, the overall score of MMedAgent is 1.8 times higher than that of LLaVA-Med. We also consider LLaVA-Med (Tool in Test), an enhanced version of LLaVA-Med that incorporates the internal output of tools. MMedAgent maintains its superior performance in this case. Furthermore, the scores for organ grounding, disease grounding, and MRG exceed 100%, indicating that MMedAgent surpasses GPT-4o in these tasks. These results underscore the superior efficiency of MMedAgent in diverse medical tasks across various modalities.
@@ 5.1.3 Case Study
5.2 Open-ended Medical Dialogue
To evaluate the capability of visual question-answering tasks, we follow the setting of open-ended medical dialogue in LLaVA-Med Li et al. (2023). Here, we use the same test data as LLaVA-Med, which consists of 193 novel questions and 50 unseen images from PMC-15M Zhang et al. (2024). This dataset contains 5 modalities and can be divided into two main classes: conversation questions and detailed description questions. We also utilize the relative score, introduced in Section 5.1.1, as the evaluation criterion. Since this is a pure language task, we select the output from GPT-4 rather than GPT-4o as the reference score.
As shown in Table 3, performance is categorized by either question types (conversation and description) or image modalities (X-ray, MRI, Histology, Gross and CT). After instruction-tuning on the tool learning dataset, MMedAgent performs better on both types of questions. Moreover, MMedAgent outperforms LLaVA-Med in all domains but MRI. This demonstrates the efficiency of MMedAgent in open-ended medical dialogue.
@@ 5.3 Tool Utilization
One significance of MMedAgent is its ability to adapt to new tools. Here, we consider two scenarios. Firstly, when a superior tool for tasks that MMedAgent is already equipped to handle becomes available, the API name of the outdated tool can be seamlessly replaced with that of the new tool, eliminating the need for additional retraining. Secondly, to extend MMedAgent to a new task, it is sufficient to generate a small set of instruction-tuning data for this specific task and fine-tune the agent accordingly, rather than retraining it from the beginning.
To verify this capability, we simulate a new tool called “Pseudo Tool”, generate an additional 5K instruction-tuning data (following Section 4), and create 30 unseen diverse questions for evaluation following Section 5.1.1. We utilize the same training settings to fine-tune MMedAgent with a smaller learning rate of 1e-6 and a batch size of 10 on one 80G A100 GPU. As shown in Figure 4, the accuracy of selecting a new tool increase to 100% within 2K steps without damaging the performance on selecting old tools.
@@ 7 Limitation
Our work is currently limited to seven tasks across five modalities. Due to the need for extensive domain knowledge, and the complexity and diversity of medical datasets involved in medical tasks, more specialized tools are emerging that should be included in our tools lists. However, the scalability of our model allows for the inclusion of more powerful tools in the future.
Additionally, more ablation studies on different backbones are necessary. Our current backbone is based on the LLaVA-Med, but recently, multiple generalist LLMs in the medical domain have been proposed, which could potentially be used to build a stronger MMedAgent.
@@ A.2 Retrieval Augmented Generation (RAG)
RAG distinguishes itself from standard report generation by its access to an external knowledge base, such as Merck Manual. We consider the following 3 common uses of RAG. The instruction-tuning data are generated based on these functionalities.
@@ A.3 Medical Grounding DINO
The datasets used to tune the medical grounding DINO is shown in Table 4.
@@ Appendix B Instruction Tuning Dataset Generation
We represent our prompts for generating an instruction tuning dataset in Figure 5.
@@ Appendix C Agent Serving
MMedAgent operates within the FastChat system , which consists of web servers that interact with users, model workers hosting the language model, and various tools. A controller coordinates the activities between the web servers and model workers. The entire system, including the 7B MMedAgent and all associated tools, can be run on an Nvidia A100 (80GB) GPU.
@@ Appendix D Evaluation Prompt
We utilize GPT-4 to assess the answers generated by MMedAgent and other models with prompts shown in Figure 6.

### v2

@@ 4 Experimental Settings
MMedAgent is initialized with LLaVA-Med 60K-IM, instruction-tuned using LoRA Hu et al. (2021) for 15 epochs, and conducted over approximately 72 hours on two 80G NVIDIA A100 GPUs. The rank of LoRA is set to 128, and the training batch size is set to 48. We employ AdamW Loshchilov and Hutter (2019) as the optimizer alongside a cosine learning rate schedule peaking at 2e-4. We generate 48K instruction-tuning data, consisting of 15K augmented VQA instruction following the method from LLaVA-Plus Liu et al. (2023b) derived from 60K inline mentions Li et al. (2023), 10K data points for detection, 3K for RAG, 5K each for segmentation, classification, MRG, and G-Seg. Data sources are shown in Table 1.
@@ 5.1.1 Evaluation Criterion
RAD-VQA SLAKE PATH-VQA PMC-VQA
Open Close Open Close Open Close Close
(Question Count) (150) (150) (150) (150) (150) (150) (300)
LLaVA-Med 28.23 61.40 39.17 52.16 12.30 54.05 27.48
MMedAgent 58.31 86.72 79.39 86.34 39.16 90.38 39.50
Table 4: Comparison of VQA tasks between MMedAgent and LLaVA-Med across different VQA benchmarks.
@@ 5.1.2 Experimental Results
As illustrated in Table 2, MMedAgent significantly outperforms all other baselines on various tasks. Note that RadFM cannot handle grounding and RAG tasks, and Flamingo-Med is not applicable for RAG because it cannot process text-only input. It is observed that the overall relative score of MMedAgent (109.48) outperforms all other state-of-the-art MLLMs by a large margin, being 1.8 times higher than that of LLaVA-Med (60.68), which is the backbone of MMedAgent. We also propose LLaVA-Med (Tool in Test), an enhanced version of LLaVA-Med that incorporates the internal output of tools and MMedAgent maintains its superior performance in this case.
Furthermore, the scores for organ grounding, disease grounding, and MRG exceed 100%, indicating that MMedAgent surpasses GPT-4o in these tasks. These results underscore the superior efficiency of MMedAgent in diverse medical tasks across various modalities.
@@ 5.1.3 Case Study
5.2 Medical VQA
When implementing VQA tasks, MMedAgent can rely on its backbone, i.e., LLaVA-Med (Sections 5.2.1) and could be further enhanced by leveraging other VQA specialists (5.2.2).
5.2.1 Open-ended Medical Dialogue
We follow the setting of open-ended medical dialogue in LLaVA-Med Li et al. (2023) and use the same test data as LLaVA-Med, consisting of 193 novel questions and 50 unseen images from PMC-15M Zhang et al. (2024a). This dataset contains 5 modalities and can be divided into two main classes: conversation questions and detailed description questions. We utilize the same relative score in Section 5.1.1 as the evaluation criterion. Since this is a pure language task, we select the output from GPT-4 rather than GPT-4o as the reference score.
As shown in Table 3, performance is evaluated from the perspective of question types (conversation and description) and image modalities (X-ray, MRI, Histology, Gross and CT). After instruction-tuning on the tool learning dataset, MMedAgent performs better on both types of questions. Moreover, MMedAgent outperforms LLaVA-Med in all domains but MRI, demonstrating the efficiency of MMedAgent in open-ended medical dialogue.
5.2.2 VQA Benchmark
MMedAgent is also evaluated on four VQA benchmarks, including VQA-RAD Lau et al. (2018), SLAKE Liu et al. (2021), PMC-VQA Zhang et al. (2023d) and PATH-VQA He et al. (2020). We construct the instruction-tuning data for each dataset and fine-tune the model with these newly added VQA tools. For each dataset, We randomly select 4K image-text pairs as the training set and 300 pairs for evaluation. Specifically, VQA-RAD, PATH-VQA, and SLAKE each contain 150 samples from the open set and 150 from the closed set, while PMC-VQA comprises 300 closed multi-choice questions. Figure 4 indicates that MMedAgent significantly outperforms LLaVA-Med across all VQA benchmarks.
@@ 5.3 Tool Utilization
One significance of MMedAgent is its ability to adapt to new tools. Here, we consider two scenarios. Firstly, when a superior tool for tasks that MMedAgent is already equipped to handle becomes available, the API name of the outdated tool can be seamlessly replaced with that of the new tool, eliminating the need for additional retraining. Secondly, to extend MMedAgent to a new task, it is sufficient to generate a small set of instruction-tuning data for this specific task and fine-tune the agent accordingly, rather than retraining it from the beginning. To verify this capability, we simulate a new tool called “Pseudo Tool”, generate an additional 5K instruction-tuning data (following Section 4), and create 30 unseen diverse questions for evaluation following Section 5.1.1. We utilize the same training settings to fine-tune MMedAgent with a smaller learning rate of 1e-6 and a batch size of 10 on one 80G A100 GPU. As shown in Figure 4, the accuracy of selecting a new tool increase to 100% within 2K steps without damaging the performance on selecting old tools.
@@ 7 Limitation
Our work is currently limited to seven tasks across five modalities. Due to the need for extensive domain knowledge, more specialized tools should be included and MMedAgent’s scalability allows for the inclusion of more powerful tools in the future. Additionally, more generalist LLMs in the medical domain could potentially serve as stronger backbone to enhance MMedAgent.
@@ A.2 Retrieval Augmented Generation (RAG)
RAG distinguishes itself from standard report generation by its access to an external knowledge base, such as Merck Manual. We consider the following three common uses of RAG. The instruction-tuning data are generated based on these functionalities.
@@ A.3 Medical Grounding DINO
The datasets used to fine-tune the medical grounding DINO is shown in Appendix Table 5.
@@ Appendix B Instruction Tuning Dataset Generation
We represent our prompts for generating an instruction tuning dataset in Appendix Figure 5.
@@ Appendix C Agent Serving
MMedAgent operates within the FastChat system , which consists of web UI that interact with users, model workers hosting the language model, and various tools. A controller coordinates the activities between the web UI and model workers. The entire system, including the 7B MMedAgent and all associated tools, can be run on an Nvidia A100 (80GB) GPU.
Figure 6: The user interface of the MMedAgent Web UI. Users can upload images and input questions in text, allowing MMedAgent to select the appropriate tool and provide comprehensive answers. The example shown demonstrates a request for segmenting organs in an abdominal CT image and generating diagnostic report for an X-ray image showing the thought progress of MMedAgent.
@@ Appendix D Evaluation Prompt
We utilize GPT-4 to assess the answers generated by MMedAgent and other models with prompts shown in Appendix Figure 7.